y varies indirectly as twice x. When x = 4, y = 1. Find y when x = 5.

Answer by stanbon(75887)

y varies indirectly as twice x.
---
y = k(2x)
---
Solve for "k" using "x = 4 when y = 1".
1 = k(8)
k = 1/8
-----
Equation:
y = (1/8)(2x) = (1/4)x
-------------------------------
Find y when x = 5.
-----
y = (1/4)5
---
y = 5/4
==============
Cheers,
Stan H.
===============